OpenStudy (anonymous):

i need to solve this rational expression: 3/x + 4/x+2 = -1/x

OpenStudy (gw2011):

I assume the problem is as follows: (3/x) + (4/x) + 2 = (-1/x) 3/x + 4/x +1/x = -2 8/x = -2 8 = -2x Rearranging this to get: -2x = 8 x = -4

OpenStudy (anonymous):

(3/x)+4/(x+2)=(-1/x)

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

{3(x+2) +4x}/ x(x+2) = -1(x+2)/x(x+2) Cancelling 1/ x(x+2), 3(x+2) + 4x = -1(x+2) 3x + 6 + 4x = -x - 2 8x = -8 x = -1

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

@shine & bbv1814 - Try substituting the values: p(-1) : LHS = 3/-1 + 4/ (-1+2) = -3 + 4 = 1 RHS = -1/-1 = 1 Therefore, LHS = RHS so the solution x = -1 is true. If you use x = -4, LHS = 3/-4 + 4/ (-4+2) = -3/4 -2 = -11/4 RHS = -1/-4 = 1/4 Clearly, LHS is not equal to RHS so x = -4 is not a solution

OpenStudy (gw2011):

Since the problem is not as I assumed orginally but has been clarified, then the solution is different which is the following: 3/x + 4/(x+2) = -1/x Multiply both sides by x and you get: 3 + 4x/(x+2) = -11 Multiply both sides by (x+2) and you get: 3(x+2) +4x = -1(x+2) 3x + 6 +4x = -x-2 7x + 6 = -x-2 8x = -8 x = -1